# arXiv:2602.24042v1[cs.DS] 27 Feb 2026

### v1

@@ 5 Power of Constant-Many Adaptive Choices: $k$-$n$ Gap
Let $I$ be a stochastic knapsack instance, and let $I_{1},I_{2}$ be a partition of $I$ into two sub-instances. Suppose $A_{1}$ and $A_{2}$ are approximation algorithms for $I_{1}$ and $I_{2}$ with approximation ratios $c_{1}$ and $c_{2}$, respectively.
Then the algorithm that partitions $I$ into $I_{1}$ and $I_{2}$ and runs either $A_{1}$ on $I_{1}$ or $A_{2}$ on $I_{2}$ (whichever yields the higher expected value) achieves an approximation ratio of at most $c_{1}+c_{2}$.
@@ 5.1 Small Items
Unfortunately, this simple approach (of inserting the items according to the greedy ordering, non-increasing order of $\frac{w_{i}}{\mu_{i}}$) fails for $\operatorname{Risky-SK}$: it has an unbounded adaptivity gap, as it may overflow w.p. $1$ (e.g. a simple instance with identical items with value $1$ and deterministic size $\varepsilon$).
@@
In this section we show a semi-adaptive policy that achieves a large fraction (approximately $\frac{1}{e}$) of $\Phi(1)$ for an instance consisting of only small items. Our algorithm, Algorithm 3, receives two parameters: $k$, the number of adaptive decisions; and $\alpha=(\alpha_{1},\ldots,\alpha_{k+1})$, a vector controlling the target expected mass of each inserted block, where $\alpha_{i}$ is the fraction of remaining capacity targeted at round $i$.
@@
Algorithm 3 orders the items according to the greedy ordering and inserts some prefix of it via inserting at most $m:=k+1$ blocks. Let $p_{j}$ denote the overflow probability of the $j$-th block $B_{j}$, conditional on the success of blocks $1$ to j1j-1. That is, $p_{j}=\Pr(Y_{j}>1\mid Y_{j-1}\leq 1)$ (since $Y_{j}$ is the used capacity of the knapsack after inserting blocks $1\ldots,j$).
@@
Let us bound $\operatorname{\mathbb{E}}[A]$ where $A$ is the total value our policy gets:
@@
    Algorithm 3 with parameters $(k=0,\alpha=0.5)$ has an adaptivity gap of at most $8$.
@@
2 highlights the trade-off between the number of adaptive choices and the quality of approximation. See Fig. 5 for illustration. Non-adaptive execution ($k=0$) yields a gap of $8$, while using $O(1/\sqrt{\varepsilon})$ adaptive choices improves the bound to $2e+\sqrt{\varepsilon}$. Remarkably, even a single adaptive decision already improves the gap from $8$ to $6.75$.
@@ 5.2 Large Items
We observe the semi-adaptive policy that follows an optimal adaptive policy for the first $k$ rounds, and then follows the $c$-approximation non-adaptive policy that we know exists for a constant $c$. For $\operatorname{Non-Risky-SK}$ we could use the $4$-approximate algorithm of [DGV08], and for $\operatorname{Risky-SK}$ we could use our $8.47$ approximate non-adaptive Algorithm 1. Choosing $c=8.5$ will work for both cases.
@@ 5.3 Implied Improvements on the Best-Known Approximation Guarantees
Our algorithm implies an improvement on the best known fully adaptive policy approximation for $\operatorname{Risky-SK}$: The $6.44+O(\sqrt{\varepsilon})$ approximation guarantee of Algorithm 3 ( 4) improves the best known $8$-approximate fully adaptive policy of [FLX18]. The improvement is both in the approximation ratio and in the number of adaptive choices.
@@ 6 Single Adaptive Choice Compared to No Adaptivity: 0-11 Gap
The class $\mathcal{H}_{k}$ may have broader applicability. In Section 7 we show that if all items are “almost” Bernoulli (taking sizes that are either close to 0 or close to $1$) then we can get adaptivity gap bounds for $\operatorname{Non-Risky-SK}$ and $\operatorname{Risky-SK}$ by analyzing the gap of $\mathcal{H}_{k}$ instances for $k>2$ and taking $k\to\infty$.
@@ 6.1 Reduction to $\mathcal{H}_{2}$
Let $I$ be a Stochastic Knapsack instance, and let $T$ be a rooted tree such that the nodes of $T$ are the items of $I$.
We say $I^{T}$ is an instance of Stochastic Knapsack with tree constraints if any policy may only insert items in $I$ that are all contained in at least one root-to-leaf path in $T$.
@@
We show that for any $\operatorname{Risky-SK}$ instance $I$, there exists an instance with tree constraints $I^{\prime T}\in\mathcal{H}_{k+1}^{T}$ such that $G_{k}(I)\leq G(I^{\prime T})$ where $G_{k}(I)$ is the $0\text{-}k$ semi-adaptivity gap of $I$, and $G(I^{\prime T})$ is the (full) adaptivity gap of $I^{\prime T}$.
Construct $I^{\prime T}$ as follows. Let $\sigma_{AD}$ be an optimal $k$-semi-adaptive policy for $I$, and let $T_{I}$ be its adaptive decision tree with $k$ adaptive adaptive decision points (where an adaptive decision point is a tree node with at least two children). Between any two adaptive decision points, the policy behaves non-adaptively—i.e., deterministically inserts a list of items $L=(i_{1},\ldots,i_{l})$.
@@
  • The adaptive policy $\sigma_{AD}$ for $I$ naturally induces a $k$-adaptive policy in $I^{\prime T}$ with the same distribution over accumulated size and value: Consider an optimal adaptive policy of instance $I$. For any root-to-leaf path in its decision tree, by replacing each list $L$ in the path by its corresponding compound item in $I^{\prime T}$, we get an adaptive policy in $I^{\prime T}$. Since the adaptive policy inserts all items between two consecutive adaptive decision points, the knapsack capacity distribution in any decision point is the same for both policies. Also, the value of items inserted between any two decision points is also the same. Thus, $\operatorname{ADAPT}(I^{\prime T})\geq\operatorname{ADAPT}(I)$.
  • Any non-adaptive policy $\sigma^{\prime}$ in $I^{\prime T}$ corresponds to one in $I$ that inserts the underlying original items in each $r_{L}$. The tree constraints make sure that for any two compound items $r_{L_{1}}$, $r_{L_{2}}$ that $\sigma^{\prime}$ attempts to insert, it must be that $L_{1}\cap L_{2}=\emptyset$, and so the resulting non-adaptive policy for $I$ is well-defined. Hence $\operatorname{ALG}(I^{\prime T})\leq\operatorname{ALG}(I)$.
@@ 6.2 Adaptivity Gap of $\mathcal{H}_{2}$
where $G(I)$ is the adaptivity gap of instance $I$. Once we narrow down to $\mathcal{H}^{\prime\prime}_{2}$, a family of instances with a specific structure, it is easier to argue about the worst adaptivity gap of an instance.
@@
The intuition is that all items in the bottom layer of the adaptive decision tree (as depicted in Fig. 7) have size bigger than half, which assures no non-adaptive policy may pick more than one of these. One the other hand, any adaptive policy picking item 0 and then some other item $i$ has the same probability of getting it as it did before this change. Determining the size distributions this way may not decrease $\operatorname{ADAPT}$ and may not increase $\operatorname{ALG}$ (the optimal adaptive decision tree is the same, but $\operatorname{\sigma_{ALG}}$ only less has options). In this case, any policy may choose to stop even after inserting a single element (to avoid overflow). Again, we get that there are only $n+1$ viable options for $\operatorname{\sigma_{ALG}}$: Either insert item 0 and then some item $i\in[n]$, or insert item $1$ and finish.
@@ 6.3 Implied Improvements on the Best-Known Approximation Guarantees
Our $\mathcal{H}_{2}$ gaps imply improvements on the best known lower bounds on the full adaptivity gaps: From $1.5$ to $1.69$ for $\operatorname{Risky-SK}$ (3) and from $1.25$ to $1+\frac{1}{e}\approx 1.37$ for $\operatorname{Non-Risky-SK}$ (4). We further improve the $\operatorname{Risky-SK}$ adaptivity gap lower bound to $2$ in Section 7 (8).
@@ 7 Limitations of a Single Adaptive Choice: 11-$n$ Gap Lower Bound
3 follows from 1 (which ties together the two notions of semi-adaptivity gaps with the full adaptivity gap). We’ve already shown a $0\text{-}1$ semi-adaptivity gap upper bound of $1+\ln(2)$ (7). We now show a (full) adaptivity gap lower bound of $2$ (8), and thus 3 follows.
@@
We show that the adaptivity gap for this family of instances is at least $2$ for $\operatorname{Risky-SK}$ (Section 7.1). For $\operatorname{Non-Risky-SK}$, we show that the adaptivity gap for this family is at most $2$ (Section E.2).
@@ 7.1 Stronger Lower Bound on the Adaptivity Gap
In this section we show a lower bound of $2$ for the adaptivity gap of $\operatorname{Risky-SK}$.
@@
For $\operatorname{Risky-SK}$ the adaptivity gap is at least $2$. This is true even when restricting to the family of $\operatorname{\varepsilon-\text{Noisy Bernoulli}}$ instances.
The theorem is an immediate result of 10 below, to which we dedicate the remainder of the section. For the first step of our $3$-step approach, we start by focusing on the simpler family of $\operatorname{\varepsilon-\text{Noisy Bernoulli}}$ distributions: Consider an instance with $\operatorname{\varepsilon-\text{Noisy Bernoulli}}$ items: For any item $i$: $S_{i}=\varepsilon_{i}\text{ w.p. }p_{i}$, and $1-\varepsilon_{i}\text{ w.p. }1-p_{i}$, for small $\varepsilon_{i}\ll 1$. Let $w_{i}$ be the value of item $i$, and let $G_{k}$ be the adaptivity gap of the instance $I_{k}$, comprised of items $\left\{1,\ldots,k\right\}$. Finally, let $V_{k}$ be the expected value the best adaptive policy gets from instance $I_{k}$.
@@ 7.2 Implied Improvements on the Best-Known Approximation Guarantees
Interestingly, for $\operatorname{Non-Risky-SK}$, similar ideas yield an improvement on the adaptivity gap upper bound rather than the lower bound. We show that for for instances with $\operatorname{\varepsilon-\text{Noisy Bernoulli}}$ items, the adaptivity gap is at most $2$ (improving the existing $4$ upper bound gap [DGV08]). We defer this analysis to Section E.2.
@@ 8 Conclusion and Future Directions
There are many avenues for further research. One interesting direction is to improve our semi-adaptivity gap results for Stochastic Knapsack. We state one concrete direction in 2. Also, the problem of pinning down the (full) adaptivity gap is still open. As we’ve shown, any further lower bound improvements for the adaptivity gap can also lead to bounds on the semi-adaptivity gaps. Separating between non-adaptive and $k$-semi-adaptive policies for constant $k$ is another future direction. Finally, we believe our notions of semi-adaptivity gaps, as well as some of our techniques (such as the adaptive decision tree analyses via the Simplify-Equalize-Optimize $3$-step approach), may be useful for studying other combinatorial stochastic optimization problems beyond Stochastic Knapsack (e.g. Stochastic Probing, Stochastic Orienteering).
@@ A.1 Proof of 1
For $\operatorname{Non-Risky-SK}$ the number of items inserted is a negative binomial random variable (the number of successful inserts until the second realization to size $1$ leading to an overflow minus the last element): $T\sim NB(2,\varepsilon)-1$ and thus $\operatorname{\mathbb{E}}[T]=\frac{2}{\varepsilon}-1$. Since the value of each item is $\varepsilon$ we get the first property.
@@
There are two phases of execution. In the first phase, the knapsack capacity is 0 and $\operatorname{\sigma_{AD}}$ keeps inserting items (as it is the optimal thing to do - there is no risk in this phase). The first phase is over when the first size $1$ realization occurs, after $T_{1}$ time and then the second phase begins, and lasts until $\operatorname{\sigma_{AD}}$ chooses to stop or an overflow has occurred. Let $T_{2}$ denote the second phase duration. So $T=T_{1}+T_{2}$, and thus from linearity of expectation:
@@
      Stop. This action leads to state stopstop w.p. $1$ and its reward is $i$.
@@
Figure 8: Second phase MDP illustration. Insert action has probability $\varepsilon$ of overflowing and probability $1-\varepsilon$ to move to a state with one more item successfully inserted into the knapsack. Stop has probability $1$ of not overflowing and reaching a finite state.
Let $V(i)$ denote the optimal expected reward when we are in non‑absorbing state $i$. Then the Bellman equation we get is: $V(i)=\max\left\{i,(1-\varepsilon)V(i+1)\right\}$.
We will show that the optimal policy for this $MDP$ is a threshold policy and find it.
@@
Assume by contradiction that for any $i\in\mathbb{N}$ the best action is to insert. The policy will overflow w.p. $1$ (as the $MDP$ is infinite and there’s a constant probability of overflow in each step), and in no step it will acquire any reward (since positive reward is only obtained via a stopping action). Thus, the policy would always have value 0. However, the policy simply always stops for any $i$, has positive expected value. We’ve reached a contradiction, and thus there must exists a state $k$ where the optimal action is to stop. W.l.o.g. let $k$ be the minimum index for which the best action in state $k$ is to stop (otherwise we could rename $k$ as the minimum index). For any $i>k$ the optimal action in state $i$ must also be to stop since the additional expected value (over stopping) that might be gained in the future is no bigger than in state $k$, the overflow probability is the same as in state $k$, and the value obtained via stopping is bigger than in state $k$.
@@
We note that given $T_{1}$ it must be that $T_{2}=max(k-T_{1},0)$: If $T_{1}\geq k$ then the optimal decision for $\operatorname{\sigma_{AD}}$ according to our $MDP$ analysis is to stop, and otherwise ($T_{1}<k$), $\operatorname{\sigma_{AD}}$ is in state $T_{1}$ in the MDP and thus (via optimality) it will try and insert until reaching state $k$ which implies at most $T_{2}=k-T_{1}$ inserts. And so:

### v2

@@ 5 Power of Constant-Many Adaptive Choices: $k$-$n$ Gap
Let II be a stochastic knapsack instance, and let $I_{1},I_{2}$ be a partition of II into two sub-instances. Suppose $A_{1}$ and $A_{2}$ are approximation algorithms for $I_{1}$ and $I_{2}$ with approximation ratios $c_{1}$ and $c_{2}$, respectively.
Then the algorithm that partitions II into $I_{1}$ and $I_{2}$ and runs either $A_{1}$ on $I_{1}$ or $A_{2}$ on $I_{2}$ (whichever yields the higher expected value) achieves an approximation ratio of at most $c_{1}+c_{2}$.
@@ 5.1 Small Items
Unfortunately, this simple approach (of inserting the items according to the greedy ordering, non-increasing order of $\frac{w_{i}}{\mu_{i}}$) fails for $\operatorname{Risky-SK}$: it has an unbounded adaptivity gap, as it may overflow w.p. 11 (e.g. a simple instance with identical items with value 11 and deterministic size $\varepsilon$).
@@
In this section we show a semi-adaptive policy that achieves a large fraction (approximately $\frac{1}{e}$) of $\Phi(1)$ for an instance consisting of only small items. Our algorithm, Algorithm 3, receives two parameters: $k$, the number of adaptive decisions; and $\alpha=(\alpha_{1},\ldots,\alpha_{k+1})$, a vector controlling the target expected mass of each inserted block, where $\alpha_{i}$ is the fraction of remaining capacity targeted at round ii.
@@
Algorithm 3 orders the items according to the greedy ordering and inserts some prefix of it via inserting at most $m:=k+1$ blocks. Let $p_{j}$ denote the overflow probability of the $j$-th block $B_{j}$, conditional on the success of blocks 11 to j1j-1. That is, $p_{j}=\Pr(Y_{j}>1\mid Y_{j-1}\leq 1)$ (since $Y_{j}$ is the used capacity of the knapsack after inserting blocks $1\ldots,j$).
@@
Let us bound $\operatorname{\mathbb{E}}[A]$ where AA is the total value our policy gets:
@@
    Algorithm 3 with parameters $(k=0,\alpha=0.5)$ has an adaptivity gap of at most 88.
@@
2 highlights the trade-off between the number of adaptive choices and the quality of approximation. See Fig. 5 for illustration. Non-adaptive execution ($k=0$) yields a gap of 88, while using $O(1/\sqrt{\varepsilon})$ adaptive choices improves the bound to $2e+\sqrt{\varepsilon}$. Remarkably, even a single adaptive decision already improves the gap from 88 to $6.75$.
@@ 5.2 Large Items
We observe the semi-adaptive policy that follows an optimal adaptive policy for the first $k$ rounds, and then follows the $c$-approximation non-adaptive policy that we know exists for a constant $c$. For $\operatorname{Non-Risky-SK}$ we could use the 44-approximate algorithm of [DGV08], and for $\operatorname{Risky-SK}$ we could use our $8.47$ approximate non-adaptive Algorithm 1. Choosing $c=8.5$ will work for both cases.
@@ 5.3 Implied Improvements on the Best-Known Approximation Guarantees
Our algorithm implies an improvement on the best known fully adaptive policy approximation for $\operatorname{Risky-SK}$: The $6.44+O(\sqrt{\varepsilon})$ approximation guarantee of Algorithm 3 ( 4) improves the best known 88-approximate fully adaptive policy of [FLX18]. The improvement is both in the approximation ratio and in the number of adaptive choices.
@@ 6 Single Adaptive Choice Compared to No Adaptivity: 0-11 Gap
The class $\mathcal{H}_{k}$ may have broader applicability. In Section 7 we show that if all items are “almost” Bernoulli (taking sizes that are either close to 0 or close to 11) then we can get adaptivity gap bounds for $\operatorname{Non-Risky-SK}$ and $\operatorname{Risky-SK}$ by analyzing the gap of $\mathcal{H}_{k}$ instances for $k>2$ and taking $k\to\infty$.
@@ 6.1 Reduction to $\mathcal{H}_{2}$
Let II be a Stochastic Knapsack instance, and let $T$ be a rooted tree such that the nodes of $T$ are the items of II.
We say $I^{T}$ is an instance of Stochastic Knapsack with tree constraints if any policy may only insert items in II that are all contained in at least one root-to-leaf path in $T$.
@@
We show that for any $\operatorname{Risky-SK}$ instance II, there exists an instance with tree constraints $I^{\prime T}\in\mathcal{H}_{k+1}^{T}$ such that $G_{k}(I)\leq G(I^{\prime T})$ where $G_{k}(I)$ is the $0\text{-}k$ semi-adaptivity gap of II, and $G(I^{\prime T})$ is the (full) adaptivity gap of $I^{\prime T}$.
Construct $I^{\prime T}$ as follows. Let $\sigma_{AD}$ be an optimal $k$-semi-adaptive policy for II, and let $T_{I}$ be its adaptive decision tree with $k$ adaptive adaptive decision points (where an adaptive decision point is a tree node with at least two children). Between any two adaptive decision points, the policy behaves non-adaptively—i.e., deterministically inserts a list of items $L=(i_{1},\ldots,i_{l})$.
@@
  • The adaptive policy $\sigma_{AD}$ for II naturally induces a $k$-adaptive policy in $I^{\prime T}$ with the same distribution over accumulated size and value: Consider an optimal adaptive policy of instance II. For any root-to-leaf path in its decision tree, by replacing each list $L$ in the path by its corresponding compound item in $I^{\prime T}$, we get an adaptive policy in $I^{\prime T}$. Since the adaptive policy inserts all items between two consecutive adaptive decision points, the knapsack capacity distribution in any decision point is the same for both policies. Also, the value of items inserted between any two decision points is also the same. Thus, $\operatorname{ADAPT}(I^{\prime T})\geq\operatorname{ADAPT}(I)$.
  • Any non-adaptive policy $\sigma^{\prime}$ in $I^{\prime T}$ corresponds to one in II that inserts the underlying original items in each $r_{L}$. The tree constraints make sure that for any two compound items $r_{L_{1}}$, $r_{L_{2}}$ that $\sigma^{\prime}$ attempts to insert, it must be that $L_{1}\cap L_{2}=\emptyset$, and so the resulting non-adaptive policy for II is well-defined. Hence $\operatorname{ALG}(I^{\prime T})\leq\operatorname{ALG}(I)$.
@@ 6.2 Adaptivity Gap of $\mathcal{H}_{2}$
where $G(I)$ is the adaptivity gap of instance II. Once we narrow down to $\mathcal{H}^{\prime\prime}_{2}$, a family of instances with a specific structure, it is easier to argue about the worst adaptivity gap of an instance.
@@
The intuition is that all items in the bottom layer of the adaptive decision tree (as depicted in Fig. 7) have size bigger than half, which assures no non-adaptive policy may pick more than one of these. One the other hand, any adaptive policy picking item 0 and then some other item $i$ has the same probability of getting it as it did before this change. Determining the size distributions this way may not decrease $\operatorname{ADAPT}$ and may not increase $\operatorname{ALG}$ (the optimal adaptive decision tree is the same, but $\operatorname{\sigma_{ALG}}$ only less has options). In this case, any policy may choose to stop even after inserting a single element (to avoid overflow). Again, we get that there are only $n+1$ viable options for $\operatorname{\sigma_{ALG}}$: Either insert item 0 and then some item $i\in[n]$, or insert item 11 and finish.
@@ 6.3 Implied Improvements on the Best-Known Approximation Guarantees
Our $\mathcal{H}_{2}$ gaps imply improvements on the best known lower bounds on the full adaptivity gaps: From $1.5$ to $1.69$ for $\operatorname{Risky-SK}$ (3) and from $1.25$ to $1+\frac{1}{e}\approx 1.37$ for $\operatorname{Non-Risky-SK}$ (4). We further improve the $\operatorname{Risky-SK}$ adaptivity gap lower bound to 22 in Section 7 (8).
@@ 7 Limitations of a Single Adaptive Choice: 11-$n$ Gap Lower Bound
3 follows from 1 (which ties together the two notions of semi-adaptivity gaps with the full adaptivity gap). We’ve already shown a $0\text{-}1$ semi-adaptivity gap upper bound of $1+\ln(2)$ (7). We now show a (full) adaptivity gap lower bound of 22 (8), and thus 3 follows.
@@
We show that the adaptivity gap for this family of instances is at least 22 for $\operatorname{Risky-SK}$ (Section 7.1). For $\operatorname{Non-Risky-SK}$, we show that the adaptivity gap for this family is at most 22 (Section E.2).
@@ 7.1 Stronger Lower Bound on the Adaptivity Gap
In this section we show a lower bound of 22 for the adaptivity gap of $\operatorname{Risky-SK}$.
@@
For $\operatorname{Risky-SK}$ the adaptivity gap is at least 22. This is true even when restricting to the family of $\operatorname{\varepsilon-\text{Noisy Bernoulli}}$ instances.
The theorem is an immediate result of 10 below, to which we dedicate the remainder of the section. For the first step of our 33-step approach, we start by focusing on the simpler family of $\operatorname{\varepsilon-\text{Noisy Bernoulli}}$ distributions: Consider an instance with $\operatorname{\varepsilon-\text{Noisy Bernoulli}}$ items: For any item $i$: $S_{i}=\varepsilon_{i}\text{ w.p. }p_{i}$, and $1-\varepsilon_{i}\text{ w.p. }1-p_{i}$, for small $\varepsilon_{i}\ll 1$. Let $w_{i}$ be the value of item $i$, and let $G_{k}$ be the adaptivity gap of the instance $I_{k}$, comprised of items $\left\{1,\ldots,k\right\}$. Finally, let $V_{k}$ be the expected value the best adaptive policy gets from instance $I_{k}$.
@@ 7.2 Implied Improvements on the Best-Known Approximation Guarantees
Interestingly, for $\operatorname{Non-Risky-SK}$, similar ideas yield an improvement on the adaptivity gap upper bound rather than the lower bound. We show that for for instances with $\operatorname{\varepsilon-\text{Noisy Bernoulli}}$ items, the adaptivity gap is at most 22 (improving the existing 44 upper bound gap [DGV08]). We defer this analysis to Section E.2.
@@ 8 Conclusion and Future Directions
There are many avenues for further research. One interesting direction is to improve our semi-adaptivity gap results for Stochastic Knapsack. We state one concrete direction in 2. Also, the problem of pinning down the (full) adaptivity gap is still open. As we’ve shown, any further lower bound improvements for the adaptivity gap can also lead to bounds on the semi-adaptivity gaps. Separating between non-adaptive and $k$-semi-adaptive policies for constant $k$ is another future direction. Finally, we believe our notions of semi-adaptivity gaps, as well as some of our techniques (such as the adaptive decision tree analyses via the Simplify-Equalize-Optimize 33-step approach), may be useful for studying other combinatorial stochastic optimization problems beyond Stochastic Knapsack (e.g. Stochastic Probing, Stochastic Orienteering).
@@ A.1 Proof of 1
For $\operatorname{Non-Risky-SK}$ the number of items inserted is a negative binomial random variable (the number of successful inserts until the second realization to size 11 leading to an overflow minus the last element): $T\sim NB(2,\varepsilon)-1$ and thus $\operatorname{\mathbb{E}}[T]=\frac{2}{\varepsilon}-1$. Since the value of each item is $\varepsilon$ we get the first property.
@@
There are two phases of execution. In the first phase, the knapsack capacity is 0 and $\operatorname{\sigma_{AD}}$ keeps inserting items (as it is the optimal thing to do - there is no risk in this phase). The first phase is over when the first size 11 realization occurs, after $T_{1}$ time and then the second phase begins, and lasts until $\operatorname{\sigma_{AD}}$ chooses to stop or an overflow has occurred. Let $T_{2}$ denote the second phase duration. So $T=T_{1}+T_{2}$, and thus from linearity of expectation:
@@
      Stop. This action leads to state stopstop w.p. 11 and its reward is ii.
@@
Figure 8: Second phase MDP illustration. Insert action has probability $\varepsilon$ of overflowing and probability $1-\varepsilon$ to move to a state with one more item successfully inserted into the knapsack. Stop has probability 11 of not overflowing and reaching a finite state.
Let $V(i)$ denote the optimal expected reward when we are in non‑absorbing state ii. Then the Bellman equation we get is: $V(i)=\max\left\{i,(1-\varepsilon)V(i+1)\right\}$.
We will show that the optimal policy for this MDPMDP is a threshold policy and find it.
@@
Assume by contradiction that for any $i\in\mathbb{N}$ the best action is to insert. The policy will overflow w.p. 11 (as the MDPMDP is infinite and there’s a constant probability of overflow in each step), and in no step it will acquire any reward (since positive reward is only obtained via a stopping action). Thus, the policy would always have value 0. However, the policy simply always stops for any $i$, has positive expected value. We’ve reached a contradiction, and thus there must exists a state $k$ where the optimal action is to stop. W.l.o.g. let $k$ be the minimum index for which the best action in state $k$ is to stop (otherwise we could rename $k$ as the minimum index). For any $i>k$ the optimal action in state $i$ must also be to stop since the additional expected value (over stopping) that might be gained in the future is no bigger than in state $k$, the overflow probability is the same as in state $k$, and the value obtained via stopping is bigger than in state $k$.
@@
We note that given $T_{1}$ it must be that $T_{2}=max(k-T_{1},0)$: If $T_{1}\geq k$ then the optimal decision for $\operatorname{\sigma_{AD}}$ according to our MDPMDP analysis is to stop, and otherwise ($T_{1}<k$), $\operatorname{\sigma_{AD}}$ is in state $T_{1}$ in the MDP and thus (via optimality) it will try and insert until reaching state $k$ which implies at most $T_{2}=k-T_{1}$ inserts. And so: